OpenStudy (iwanttogotostanford):

HELP PLEASE ASAP URGENT: A light bulb whose resistance is 240 ohms is connected to a 120 voltage source. What is the current through the bulb? Identify all of your data, write the equation, and show all work.

OpenStudy (palwandew):

Do you know the formula to find current?

OpenStudy (iwanttogotostanford):

this? P = I × V. P = power

OpenStudy (palwandew):

Nope V=IR I=V/R. Now put the given values.

OpenStudy (palwandew):

I=current V=voltage R=resistance

OpenStudy (iwanttogotostanford):

ok thank you ill put those in really quick

OpenStudy (iwanttogotostanford):

so it would be I=120/240

OpenStudy (palwandew):

Yep.

OpenStudy (iwanttogotostanford):

and then what do i do from there to complete what the question is asking?

OpenStudy (iwanttogotostanford):

so the current would be 0.5 correct?

OpenStudy (palwandew):

Yes and don't forget to write its unit.

OpenStudy (iwanttogotostanford):

ok and then is that all i had to do for this question??

OpenStudy (palwandew):

Yeah thats it.

OpenStudy (iwanttogotostanford):

thanks so much!